If the complement of an angle is 6° less than one-third of its supplement, find the measure of the angle.
Answer by mananth(16946)

let angle be x
complement = 90-x
supplement = 180-x
....
90-x = 1/3 (180-x)-6
3(90-x)=180-x
270-3x=180-x
270-180+6=3x-x
96=2x
x=48